# An Extensible Software Transport Layer for GPU Networking

### v1

@@ 5. Implementation
We implement UCCL in 27.4K lines of C++ as a network plugin to NCCL, the standard collective library for NVIDIA GPUs. Our current implementation supports NVIDIA RDMA NICs via RC and UC, AWS EFA NICs via UD, and AWS ENA non-RDMA NICs via AF_XDP. UCCL leverages the popular libibverbs for RDMA NICs and kernel built-in AF_XDP for non-RDMA NICs; therefore, it should naturally support other vendors’ NICs. To enable scattered memcpy when using UD, we add or modify 170 LoC to the NCCL codebase, including a scattered memcpy GPU kernel and C++ code that passes packet pointers to the kernel via CPU-GPU shared memory (created by cudaHostAlloc()). We added two new interfaces to the NCCL network plugin system: irecv_scattered that receives data in a list of scattered packets, and irecv_free_ptrs that frees the scattered packet buffers after scattered memcpy. UCCL also supports running NCCL with multiple processes.
Name | # of Servers | Network | Topology | NIC | GPU | CPU |
CX | 2 | InfiniBand | Same rack | NVIDIA ConnectX-7 400G 8 | H100-80G 8 | Intel Xeon Platinum 8462Y+ (128 cores) |
EFA | 4 | Ethernet | Fat-tree cross racks | AWS EFA 100G 4 | A100-40G 8 | Intel Xeon Platinum 8575CL (96 cores) |
@@ 6. Evaluation
Testbed. Table 2 describes our evaluation testbeds. The CX testbed under the same rack mainly serves as a performance stress-testing for UCCL software implementation, as it has no network congestion or flow collisions. We failed to rent a larger testbed from public cloud providers because they have strict GPU quota limits for individuals. Instead, we try to simulate a larger testbed by disabling NCCL NVLink and SHM communication for some tests. This forces GPUs within a server to communicate through the network, and each GPU essentially behaves like a virtual server, e.g., CX becomes 16 virtual GPU servers each with 400G bandwidth.
Experiment setup. For collective performance evaluation, we use NCCL v2.23.4 (Corporation, 2024b) and NCCL-tests 9d26b84 (Corporation, 2024a), the latest versions when we start integrating UCCL into NCCL, and focus on the representative allreduce and all-to-all collectives. NCCL-tests vary the collective message size and measure the achieved latency and bus bandwidth; the latency is inversely proportional to the bandwidth.
Comparison baselines. On the CX testbed, we compare UCCL to the NCCL built-in RDMA support that uses CX-7 RC. UCCL uses 256 UC/RC QPs for each NIC pair, while the NCCL built-in uses QP scaling with 4 RC QPs per connection (Gangidi et al., 2024; Le et al., 2024). UCCL uses CUBIC CC on CX unless specified, which performs better than Swift in the lossless InfiniBand. Since InfiniBand has no packet loss for CUBIC to reduce cwnd, to avoid severe network congestion, we enforce a maximum cwnd value in UCCL CUBIC CC to limit the maximum inflight bytes, similar to TCP flow control.
@@
As the cost of extensible software transport, NCCL atop UCCL uses more CPU cores than vanilla NCCL. By default, the vanilla NCCL uses 2 cores per GPU, while UCCL uses 2 more cores per NIC to run the engine threads (3 more cores per NIC for receiver-driven CC due to the pacer). For example, on CX, they use 16 and 32 cores respectively; on EFA, 16 and 24 respectively.
@@ 6.1. Collective Performance
On CX. Figure 5 compares the allreduce and all-to-all performance of UCCL vs. ConnextX-7 on the CX testbed. UCCL UC/RC performs almost the same as ConnectX-7 with various data sizes, except that UCCL UC performs slightly worse (i.e., less than 4%) than ConnextX-7 and UCCL RC for allreduce after 128MB. For allreduce, this exception is because of the extra overhead of handling packet reliability in UC; for all-to-all, ConnextX-7 and UCCL RC do not perform better than UCCL UC, because all-to-all is connection-intensive, causing more QP swapping for RC. UCCL exhibits lower latency for small messages than ConnectX-7 since UCCL adaptively bypasses multipath (§3.3), while ConnectX-7 always uses 4 RC QPs per connection. These experiments confirm that UCCL can make highly efficient control decisions in software, reaching the performance of ASIC-based RDMA NICs (i.e., NVIDIA ConnectX-7) for ML collectives.
@@
On EFA. Figure 6 compares the collective performance of UCCL and SRD on the EFA testbed. UCCL CUBIC and EQDS achieve similar performance with significant improvement over SRD (except for 256MB allreduce): for allreduce, UCCL outperforms SRD by up to 1.37 and 1.27 for small and large message ranges; for all-to-all, the speedup is up to 2.93 and 3.27 respectively. UCCL outperforms SRD because beefy CPU cores are faster in making transport decisions than wimpy ARM cores on EFA NICs (that run SRD), especially when handling connection-intensive all-to-all. Surprisingly, simple CUBIC CC performs very well. This is mainly because UCCL leverages hundreds of network paths and avoids congested ones before sending packets; so most of the time, CC does not get involved. The small performance gap after 256MB is caused by the extra control header added by UCCL, while SRD directly reuses the UDP header as its control header (UCCL cannot reuse as it is not exposed to CPU software).
Figure 7 shows the collective performance comparison on the EFA testbed when NVLink and SHM are enabled (meaning a smaller testbed). Even though a major amount of data traffic goes through the high-bandwidth NVLink, UCCL still achieves much higher or comparable collective performance than SRD, i.e., up to 1.57 for allreduce and 2.14 for all-to-all. We note that all-to-all with NVLink on performs worse than NVLink off (Figure 6). This is because our current UCCL implementation does not support the NCCL PXN (NVIDIA Corporation, 2022) optimization when NVLink is on. Basically, PXN aggregates network messages destined to the same server, and uses the destination NVLink to segment them into different GPUs, which vastly improves small message latency. However, it involves a complicated multireceive mechanism that causes engineering difficulties for UCCL UD integration (that requires changing the NCCL codebase), which we leave as future development. Nevertheless, UCCL still performs better than SRD under small messages in this experiment.
@@ 6.2. Application Performance
We run two applications to evaluate how UCCL improves ML workload performance on the EFA testbed. One is a classic ResNet (He et al., 2016) distributed training in PyTorch with the widely-used 1F1B mechanism (Harlap et al., 2018; Zhang et al., 2017) that tries to overlap communication with computation. However, as the GPU compute capability grows much faster than the network bandwidth, the communication time cannot be entirely hidden by the computation time, especially on EFA. Another is a DeepSeek-V3-like MoE serving application. However, we fail to directly serve DeepSeek-V3 on our testbed, because (1) it uses a customized communication library rather than the standard NCCL, so it cannot directly use UCCL, and (2) our testbed does not have enough GPU memory to fit all model weights. Instead, we get a realistic DeepSeek-V3 trace including GPU compute time and network message sizes (i.e., hidden state sizes) from (zartbot/shallowsim, 2025), and emulate the computation and communication behaviors with PyTorch and NCCL.
@@ 6.3.1. Handling network incast
In a shared RDMA network, network incast can lead to victim flows due to PFC propagation (Wang et al., 2023; Hu et al., 2016; Mittal et al., 2018). Such network incast can be caused by the Expert Parallelism in MoE serving (§2.2) and gather/reduce collectives in multi-level parallelism of LLM training (Barham et al., 2022; Grattafiori et al., 2024). This experiment tries to create a scenario where network incast co-exists with other collective traffic, and evaluates how UCCL performs compared to the RDMA hardware transport on CX. We co-locate 15-to-1 incast traffic and 16-NIC permutation traffic (Raiciu et al., 2011; Cao et al., 2013; Le et al., 2024; Bonato et al., 2024), where each NIC streams data to another and no NIC receives more than one stream, representing typical collectives. For both types of traffic, each NIC sends 1MB messages with at most four in flight. We compare the receiver-driven EQDS in UCCL vs. the sender-driven InfiniBand CC in ConnectX-7.
Figure 9 and 9 show the FCT distributions for incast traffic and permutation traffic, respectively. Compared to InfiniBand, UCCL EQDS reduces P99/P99.9 latency by 1.73/1.72 for incast traffic and 4.50/4.88 for permutation traffic. This is because the InfiniBand CC only reduces the sending rate when severe congestion and queue build-up have occurred on the incast switch port; but this is too late, as the Credit-Based Flow Control (InfiniBand Trade Association, 2020) (PFC equivalent in InfiniBand) has already paused all upstream ports and heavily disturbed victim flows (i.e., permutation traffic). Conversely, UCCL EQDS proactively controls the rate of all senders on the receiver side, which reduces queue build-up and avoids upstream ports entering the pausing state.
@@ 6.3.2. Handling packet loss
We then evaluate how UCCL selective retransmission handles packet loss, compared to hardware-based loss recovery. Furthermore, UCCL coalesces transport decisions like loss recovery at chunk granularity for high software efficiency; therefore, we also test whether this design choice would cause poor loss recovery performance. To this end, we instrument different packet loss rates in software when running NCCL collectives over UCCL UC. Unfortunately, we cannot instrument packet loss for our CX NICs as that would require reconfiguring the NIC or switch behaviors, both of which we do not have access to. Instead, we cite comparable numbers from prior literature.
Figure 10 shows the results. UCCL has only 3-23%/22-41% performance drop for allreduce and all-to-all under 0.01% packet loss rate, compared to the reported 77% drop of RDMA hardware transport in Flor (Li et al., 2023) (combining Flor Figure 6(a) and 7(a) under 1/16384 loss rate). With 1% high packet loss rate, UCCL still retains 58%/72% of the performance for 1GB messages, compared to old generation RDMA NICs that achieve near-zero throughput as reported in DCQCN (Zhu et al., 2015) (Figure 21). This confirms the effectiveness and advantage of UCCL loss recovery despite control coalescing.
@@ 6.4. UCCL Scalabality
6.4.1. Varying number of QPs
We vary the number of QPs per NIC pair from 256 to 4096; given that 1 NIC connects to 15 other NICs in all-to-all, this results in 3840 (3.75K) to 61440 (60K) QPs per NIC. With 60K QPs per NIC, we essentially model the QP swapping overhead for UCCL all-to-all across 241 GPUs using 256 paths per connection (i.e., 61440/256=240, plus 1 to include itself) or 961 GPUs using 64 paths. This has covered the largest-scale collective as far as we know, i.e., 256 GPUs reported by Meta (Gangidi et al., 2024) (DeepSeek-V3 limits each token to be dispatched to 4 servers, though with 320 GPUs in total (Liu et al., 2024)).
Figure 11 shows the results. For RC, using 60K QPs incurs a maximum 17% performance drop compared to 3.75K QPs (that have similar performance as the vanilla NCCL). For UC, the performance remains nearly unchanged across different numbers of QPs. As mentioned in §3.2, this is because UCCL’s chunked data transfers (i.e., control coalescing) for large ML messages effectively amortize the QP swapping overhead while GPUDirect avoids PCIe traffic contention. Note that UCCL on EFA does not have QP scalability issues, as UCCL uses connection-less UD there.
@@ 6.4.2. Varying chunk size
Figure 12 shows the all-to-all performance on the CX testbed when varying chunk sizes. Small message performance mostly remains unaffected by different chunk sizes, thus omitted here. Our results reveal that large messages do require a larger chunk size to achieve line rate, i.e., 16KB or 4 MTU-sized chunks. This experiment demonstrates the performance benefits of control coalescing for UCCL software transport.
@@ 6.4.3. Varying number of CPU cores
Figure 12 shows the all-to-all performance when varying the number of UCCL engines/CPU cores per NIC. Small message latency exhibits negligible difference across different CPU cores, thus omitted here. For ASIC-based NICs with segmentation and reassembly offloading, even though one CPU core can only saturate bidirectional 29GB/s, two CPU cores are enough to saturate the full 50GB/s line rate. This experiment confirms the high efficiency of UCCL software transport, i.e., 1 CPU core saturating 400G unidirectional traffic.
@@ 6.5.1. Connection splitting
We now evaluate how connection splitting impacts collective performance in UCCL. Without connection splitting, each UCCL connection is mapped to a single engine out of the two UCCL engines per NIC; with splitting, each UCCL connection dispatches and load balances its messages among the two UCCL engines. Our experiment results show that UCCL without connection splitting only achieves 45.7 and 39.9 GB/s max busbw for allreduce and all-to-all on CX respectively, while UCCL with splitting achieves 48.9 and 48.5 GB/s respectively, saturating the line rate. This is because connection splitting enables multiple CPU cores to process inflight messages in a load-balanced manner. We anticipate connecting splitting will become more critical as the NIC scales to 800 Gbps (Corporation, 2025c).
6.5.2. Congestion signal fidelity in software

### v2

@@ 4.3. Efficient Loss Recovery
Name | # of Servers | Network | Topology | MTU | NIC | GPU | CPU |
CX_IB | 2 | InfiniBand | Same rack | 4KB | NVIDIA ConnectX-7 400G 8 | NVIDIA H100-80G 8 | 128 cores |
CX_ETH | 6 | Ethernet | Cross racks, fat-tree | 4KB | NVIDIA ConnectX-7 400G 8 | NVIDIA H100-80G 8 | 160 cores |
AMD | 4 | Ethernet | Cross racks, rail-optimized | 4KB | Broadcom Thor‑2 400G 7 | AMD MI300X-192G 7 | 128 cores |
EFA | 4 | Ethernet | Cross racks, fat-tree | 9KB | AWS EFA 100G 4 | NVIDIA A100-40G 8 | 96 cores |
@@ 5. Implementation
We implement UCCL in 28.4K lines of C++ as a network plugin to NCCL/RCCL, the standard collective library for NVIDIA/AMD GPUs. Our current implementation supports NVIDIA RDMA NICs via RC and UC, Broadcom RDMA NICs via RC, AWS EFA NICs via UD/SRD, and AWS ENA non-RDMA NICs via AF_XDP. UCCL leverages the standard libibverbs for RDMA NICs and kernel built-in AF_XDP for non-RDMA NICs; therefore, it should naturally support other vendors’ NICs. To enable scattered memcpy when using UD for AWS EFA NICs, we add or modify 170 LoC to the NCCL codebase, including a scattered memcpy GPU kernel and C++ code that passes packet pointers to the kernel via CPU-GPU shared memory (created by cudaHostAlloc()). We added two new interfaces to the NCCL network plugin system: irecv_scattered that receives data in a list of scattered packets, and irecv_free_ptrs that frees the scattered packet buffers after scattered memcpy. UCCL also supports running NCCL with multiple processes.
@@ 6. Evaluation
Testbed. Table 1 describes our evaluation testbeds. The CX_IB testbed under the same rack mainly serves as a performance stress-testing for UCCL software implementation, as it has no network congestion or flow collisions. The CX_ETH testbed across two racks evaluates how UCCL handles network congestion and flow collisions. The AMD and EFA testbed across racks evaluate the genericity of UCCL (i.e., applying to AMD GPUs and Broadcom/EFA NICs). Each AMD server has one NIC down out of 8; therefore, we only use 7 GPUs for evaluation. For some tests, we try to simulate a larger testbed by disabling NCCL NVLink (or RCCL xGMI) and SHM communication. This forces GPUs within a server to communicate through the network, and each GPU essentially behaves like a virtual server, e.g., CX_IB becomes 16 virtual GPU servers each with 400G bandwidth.
Experiment setup. For collective performance evaluation, we use NCCL v2.23.4 (Corporation, 2025d) and NCCL-tests 9d26b84 (Corporation, 2025c), the latest versions when we start integrating UCCL into NCCL, and focus on the representative allreduce and all-to-all collectives. NCCL-tests vary the collective message size and measure the achieved bus bandwidth. We focus on message sizes from 1MB to 1GB, commonly used in real ML workloads (Qian et al., 2024; Wang et al., 2025; Xu et al., 2025). For AMD GPUs, we use RCCL 532f54c (Advanced Micro Devices, 2025b) and RCCL-test 5b27b96 (Advanced Micro Devices, 2025a).
Comparison baselines. On the CX_IB testbed, we compare UCCL to the NCCL built-in RDMA support that uses CX-7 RC. UCCL uses 256 UC/RC QPs for each NIC pair, while the NCCL built-in uses QP scaling with 4 RC QPs per connection (Gangidi et al., 2024; Le et al., 2024). UCCL uses CUBIC CC on CX_IB unless specified, which performs better than Swift in the lossless InfiniBand. Since InfiniBand has no packet loss for CUBIC to reduce cwnd, to avoid severe network congestion, we enforce a maximum cwnd value in UCCL CUBIC CC to limit the maximum inflight bytes, similar to TCP flow control.
@@
As the cost of extensible software transport, NCCL atop UCCL uses more CPU cores than vanilla NCCL. By default, the vanilla NCCL uses 2 cores per GPU, while UCCL uses 2 more cores per NIC to run the engine threads, with 1 additional core per NIC for receiver-driven CC due to the pacer.
@@ 6.1. Collective Performance
On CX_IB. Figure 7 compares the allreduce and all-to-all performance of UCCL vs. ConnextX-7 on the CX_IB testbed. UCCL UC/RC performs almost the same as ConnectX-7 with various data sizes, except that UCCL UC performs slightly worse (i.e., less than 4%) than ConnextX-7 and UCCL RC for allreduce after 128MB. For allreduce, this exception is because of the extra overhead of handling packet reliability in UC; for all-to-all, ConnextX-7 and UCCL RC do not perform better than UCCL UC, because all-to-all is connection-intensive, causing more QP swapping for RC. These experiments confirm that UCCL can make highly efficient control decisions in software, reaching the performance of ASIC-based RDMA NICs (i.e., NVIDIA ConnectX-7) for ML collectives.
On CX_ETH. Figure 8 compares the performance of UCCL vs. ConnextX-7 on the CX_ETH testbed. For NCCL over CX-7, we use multiple QPs for each connection between two GPUs by adjusting NCCL_IB_QPS_PER_CONNECTION. However, it still suffers from significant performance drops under larger message sizes and lower peak performance. This is because flow collisions in the fat-tree topology cause severe network congestion, which in turn leads to exponential backoffs in the CX-7 CC mechanism. Instead, UCCL is able to scale up performance stably with message size increasing, by making smarter LB and CC decisions in software. Overall, UCCL outperforms CX-7 by up to 2.32/1.60/1.24 (corresponding to 4/8/16 QPs) and 1.79/3.82/4.54 for allreduce and all-to-all, respectively.
On AMD. Figure 9 shows the performance comparison on the AMD testbed. This testbed has a rail-optimized topology that helps reduce network congestion when used together with the PXN technique (NVIDIA Corporation, 2022). Even under this topology, UCCL still outperforms Thor-2 by up to 1.34/1.23/1.08 (corresponding to 4/8/16 QPs) and 1.62/1.62/1.93 for allreduce and all-to-all, by better controlling network congestion.
@@
On EFA. Figure 10 compares the collective performance of UCCL and SRD on the EFA testbed. UCCL CUBIC and EQDS achieve similar performance with significant improvement over SRD (except for 256MB allreduce): for allreduce, UCCL outperforms SRD by up to 1.27; for all-to-all, the speedup is up to 3.27. UCCL outperforms SRD because beefy CPU cores are faster in making transport decisions than wimpy ARM cores on p4d.24xlarge EFA NICs (that run SRD), especially when handling connection-intensive all-to-all. Surprisingly, simple CUBIC CC performs very well. This is mainly because UCCL leverages hundreds of network paths and avoids congested ones before sending packets; so most of the time, CC does not get involved. The small performance gap after 256MB is caused by the extra control header added by UCCL, while SRD directly reuses the UDP header as its control header (UCCL cannot reuse as it is not exposed to CPU software).
UCCL could also directly leverage the SRD protocol of EFA NICs to send/recv individual datagram packets, denoted as UCCL SRD. UCCL SRD does connection management and multipath load balancing in CPU, and packet reassembly in GPU (§3.1), overcoming the limited processing power and cache capacity of SmartNICs on p4d.24xlarge. As a result, it achieves similar high performance to UCCL CUBIC and EQDS.
Figure 11 shows the collective performance comparison on the EFA testbed when NVLink and SHM are enabled (meaning a smaller testbed). Even though a major amount of data traffic goes through the high-bandwidth NVLink, UCCL still achieves much higher or comparable collective performance than SRD, i.e., up to 1.57 for allreduce and 2.14 for all-to-all.
@@ 6.2. Application Performance
UCCL is motivated by the practical needs of both LLM training and serving workloads. We run two applications to evaluate how UCCL improves ML workload performance on the EFA testbed. One is a classic ResNet (He et al., 2016) distributed training in PyTorch with the widely-used 1F1B mechanism (Harlap et al., 2018; Zhang et al., 2017) that tries to overlap communication with computation. However, as the GPU compute capability grows much faster than the network bandwidth, the communication time cannot be entirely hidden by the computation time, especially on EFA. Another is a DeepSeek-V3-like MoE serving application. However, directly serving DeepSeek-V3 on our testbed is impractical, because (1) it uses a customized communication library rather than the standard NCCL, so it cannot directly use UCCL, and (2) it requires a large number of GPUs to run meaningful expert parallelism, e.g., 320 GPUs by DeepSeek (Liu et al., 2024) and 96 GPUs by SGLang (The SGLang Team, 2025). Instead, we get a realistic DeepSeek-V3 trace including GPU compute time and network message sizes (i.e., hidden state sizes) from (zartbot/shallowsim, 2025), and emulate the computation and communication behaviors with PyTorch and NCCL.
@@ 6.3.1. Handling network incast
In a shared RDMA network, network incast can lead to victim flows due to PFC propagation (Wang et al., 2023; Hu et al., 2016; Mittal et al., 2018). Such network incast can be caused by the Expert Parallelism in MoE serving (§2.2) and gather/reduce collectives in multi-level parallelism of LLM training (Barham et al., 2022; Grattafiori et al., 2024). This experiment tries to create a scenario where network incast co-exists with other collective traffic, and evaluates how UCCL performs compared to the RDMA hardware transport on CX_IB. We co-locate 15-to-1 incast traffic and 16-NIC permutation traffic (Raiciu et al., 2011; Cao et al., 2013; Le et al., 2024; Bonato et al., 2024), where each NIC streams data to another and no NIC receives more than one stream, representing typical collectives. For both types of traffic, each NIC sends 1MB messages with at most four in flight. We compare the receiver-driven EQDS in UCCL vs. the sender-driven InfiniBand CC in ConnectX-7.
Figure 13 and 13 show the FCT distributions for incast traffic and permutation traffic, respectively. Compared to InfiniBand, UCCL EQDS reduces P99/P99.9 latency by 1.73/1.72 for incast traffic and 4.50/4.88 for permutation traffic. This is because the InfiniBand CC only reduces the sending rate when severe congestion and queue build-up have occurred on the incast switch port; but this is too late, as the Credit-Based Flow Control (InfiniBand Trade Association, 2020) (PFC equivalent in InfiniBand) has already paused all upstream ports and heavily disturbed victim flows (i.e., permutation traffic). Conversely, UCCL EQDS proactively controls the rate of all senders on the receiver side, which reduces queue build-up and avoids upstream ports entering the pausing state.
@@ 6.3.2. Handling packet loss
We then evaluate how UCCL selective retransmission handles packet loss, compared to hardware-based loss recovery. Furthermore, UCCL coalesces transport decisions like loss recovery at chunk granularity for high software efficiency; therefore, we also test whether this design choice would cause poor loss recovery performance. To this end, we instrument different packet loss rates in software when running NCCL collectives over UCCL UC. Our packet loss instrumentation has considered the chunking nature in UCCL, where any packet loss will cause the whole chunk loss. Unfortunately, we cannot instrument packet loss for our CX_IB NICs as that would require reconfiguring the NIC or switch behaviors, both of which we do not have access to. Instead, we cite comparable numbers from prior literature Flor (Li et al., 2023).
We pick two GPUs from the AMD testbed, each using one NIC for communication. The client NIC establishes a connection with the server NIC and keeps 16 inflight messages. Similar to Flor (Li et al., 2023), we use a single QP for the connection, and disable congestion control to avoid sending rate backoff. We vary the message size (32KB1MB) and measure the goodput under different packet drop ratios. Figure 14 shows the results. UCCL has only 1% performance drop under 1/16384, 1/4096 drop ratio, compared to the reported 26%42% drop of RDMA hardware transport in Flor (Li et al., 2023, Figure 7). Even under a high drop ratio of 1/1024 and 1/256, UCCL only experiences a performance drop of 6%30%, compared to the reported 59%76% drop in RDMA hardware transport.
@@ 6.4.1. Varying chunk size
Figure 15 shows the all-to-all performance on the CX_IB testbed when varying chunk sizes. Our results reveal that saturating the line rate requires a somewhat medium chunk size, i.e., 16KB or 4 MTU-sized chunks. This experiment demonstrates the performance benefits of control coalescing for UCCL software transport.
@@ 6.4.2. Varying number of CPU cores
Figure 15 shows the all-to-all performance when varying the number of UCCL engines/CPU cores per NIC. For ASIC-based NICs with segmentation and reassembly offloading, even though one CPU core can only saturate bidirectional 29GB/s, two CPU cores are enough to saturate the full 50GB/s line rate. The slight performance drop for 16MB when switching from “2CPUs” to “4CPUs” should be caused by runtime variability. This experiment confirms the high efficiency of UCCL software transport, i.e., 1 CPU core saturating 400G unidirectional traffic (2 CPU cores for 400G bidirectional).
@@ 6.5.1. Connection splitting
We now evaluate how connection splitting impacts collective performance in UCCL. Without connection splitting, each UCCL connection is mapped to a single engine out of the two UCCL engines per NIC; with splitting, each UCCL connection dispatches and load balances its messages among the two UCCL engines. Our experiment results show that UCCL without connection splitting only achieves 45.7 and 39.9 GB/s max busbw for allreduce and all-to-all on CX_IB respectively, while UCCL with splitting achieves 48.9 and 48.5 GB/s respectively, saturating the line rate. This is because connection splitting enables multiple CPU cores to process inflight messages in a load-balanced manner. We anticipate connecting splitting will become more critical as the NIC scales to 800 Gbps (Corporation, 2025e).